# Chinese Academy of Sciences, Beijing, China

### v2

@@ 1 Introduction
A function is one-way if it is computationally easy to compute but hard to invert, even probabilistically. Many aspects of computer science such as computational complexity, pseudorandom generators and digital signatures rely on one-way functions from strings (finite binary sequences) to strings [13]. Although their existence is not known, much research effort has focused on their implications and relations with other fundamental problems in computation [10, 9, 7]. Recent research has established strong connections between one-way functions and Kolmogorov complexity [19, 14, 15, 8].
Levin [2] asked for the existence of a one-way function from reals (infinite binary sequences) to reals, which he defined in terms of computability and randomized computations. These are partial computable functions $f$ in the standard sense of computable analysis [6] which preserve algorithmic randomness and for each randomized oracle Turing machine $M$ the probability that $M$ inverts $f$, namely $f(M(y))=y$, is zero.
We give a strong positive answer to Levin’s question:
@@
There exists a total surjective one-way function $f$.
The hardness of inverting $f$ is based on the hardness of the halting problem $\emptyset^{\prime}$, which is used in the definition of $f$ via an effective enumeration of it. This allows to establish a stronger property: every randomized continuous inversion of $f$ computes $\emptyset^{\prime}$.
We also show that one-way functions are highly non-injective.
Theorem.
The inverse images $f^{-1}(y)$ of one-way functions $f$ are uncountable perfect sets almost everywhere in their range.
Finally we construct ‘nearly injective’ functions which are hard to invert deterministically or almost surely.
Independently Gács [5] constructed $f:\subseteq 2^{\omega}\to 2^{\omega}$ which is ‘one-way’ with respect to the domain of $f$ instead of its range:
  • $f$ is partial computable with domain of positive uniform measure
  • $\mu\big{(}{\{\hskip 0.85355pt{(x,r)}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0$ for each partial computable $g$.
This is a weaker condition since it is met by (a) some functions that do not map to any randoms; (b) all total random-preserving one-way functions.
@@ 2 Preliminaries
Let $\mathbb{N}$ be the set of natural numbers, represented by $n,m,i,j,t,s$. Let $2^{\omega}$ be the set of reals, represented by variables $x,y,z,v,w$, and $2^{<\omega}$ the set of strings which we represent by $\sigma,\tau,\rho$. We index the bits $x(i)$ of $x$ starting from $i=0$. The prefix of $x$ of length $n$ is $x(0)x(1)\cdots x(n-1)$ and is denoted by $x\upharpoonright_{n}$. Let $\preceq$, $\prec$ denote the prefix and strict prefix relation between two strings or a string and a real. Similarly $\succeq$,$\succ$ , ≻ denote the suffix relations and $\sigma\mid\tau$ denote $\sigma\not\preceq\tau\ \wedge\ \tau\not\preceq\sigma$. Given $x,y$ let
@@
which we call neighborhoods. Let $\mu$ be the uniform measure on $2^{\omega}$, determined by $\mu(\llbracket\sigma\rrbracket)=2^{-|\sigma|}$. Probability in $2^{\omega}\times 2^{\omega}$ is reduced to $2^{\omega}$ via the measure-preserving $(x,y)\mapsto x\oplus y$. Classes of $\mu$-measure 0 are called null.
A tree is a downward $\preceq$-closed $T\subseteq 2^{<\omega}$, is pruned if every $\sigma\in T$ has an extension in $T$ and is perfect if every $\sigma\in T$ has two extensions $\tau|\rho$ in $T$. A real is a path through $T$ if all of its prefixes belong to $T$.
Computability. We use the standard notion of relative computability in terms of Turing machines with oracles from $2^{\omega}$. Turing reducibility $x\leq_{T}z$ means that $x$ is computable from $z$ (is $z$-computable) and is a preorder calibrating $2^{\omega}$ according to computational power in the Turing degrees.
Effectively open sets are of the form $\bigcup_{i}\llbracket\sigma_{i}\rrbracket$ where $(\sigma_{i})$ is computable, and are also known as $\Sigma^{0}_{1}$ classes. Effectively closed sets, also known as $\Pi^{0}_{1}$ classes, are the complements of $\Sigma^{0}_{1}$ classes and are representable as the set of paths through a computable tree. The paths through a computable pruned tree form a $\Delta^{0}_{1}$ or decidable class. If $(\sigma_{n,i})$ is computable then
@@
A Martin-Löf test is a uniformly $\Sigma^{0}_{1}$ sequence $(V_{n})$ such that $\mu(V_{n})\leq 2^{-n}$. A real $x$ is random if $x\notin\bigcap_{n}V_{n}$ for any Martin-Löf test $(V_{n})$. A real is weakly $n$-random if it is in every $\Sigma^{0}_{n}$ set of measure $1$. Relativizations to oracle $r$ define $\Sigma^{0}_{1}(r)$ classes, weakly $r$-random reals and so on.
Computable analysis. A Turing machine with a one-way infinite output tape and access to an oracle from $2^{\omega}$ may eventually print an infinite binary sequence, hence defining a partial map from reals to reals. This standard notion of computability of real functions [17] is almost as old as computability itself [6] and implies that computable real functions are continuous. Let
@@
  • $f(x)\downarrow,f(x)\uparrow$ denote that $f(x)$ is defined or undefined.
Let $f(x;n)$ denote bit $n$ of $f(x)$ and
$f(2^{\omega}):=\{\hskip 0.85355pt{y}:{\exists x,\ f(x)=y}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf^{-1}(y):=\{\hskip 0.85355pt{x}:{f(x)=y}\hskip 0.85355pt\}$
denote the range and the inverse images of $f$. If $f$ is total, the sets
@@
are closed if $f$ is continuous and effectively closed if $f$ is computable.
Definition 2.1.
We say $\hat{f}$ is a representation of $f:\subseteq 2^{\omega}\to 2^{\omega}$ if
  • $\hat{f}:2^{<\omega}\to 2^{<\omega}$  and $\hat{f}(\lambda)=\lambda$ and $\big{(}{\sigma\preceq\tau\ \Rightarrow\ \hat{f}(\sigma)\preceq\hat{f}(\tau)}\big{)}$
  • $f(x)\downarrow\iff\lim_{\tau\prec x}\hat{f}(\tau)=f(x)\iff\lim_{\tau\prec x}|{\hat{f}(\tau)}|=\infty$
where $\lim_{\tau\prec x}\hat{f}(\tau)=z$ means $\big{(}{\forall\tau\prec x,\ \hat{f}(\tau)\prec z\ \wedge\ \ \lim_{\tau\prec x}|{\hat{f}(\tau)}|=\infty}\big{)}$.
Every continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ has a representation $\hat{f}$. If $f$ is partial computable then it has a computable representation $\hat{f}$. We fix an effective coding of the graphs of representations $\hat{f}$ into $2^{\omega}$ and identify representations with their real codes.
There is no canonical way to assign a representation to each continuous $f$: by [16] there are continuous $f$ such that for each representation $z$ of $f$ there exists another one of lesser Turing degree. So it is not possible to classify the computational power of continuous functions in the Turing degrees. This will not be an issue in our results as we only need:
@@
3 Deterministic inversions
Given continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ we say that $g:\subseteq 2^{\omega}\to 2^{\omega}$:
  • inverts $f$ on $y$ if $g(y)\downarrow$ and $f(g(y))\downarrow=y$
Injective total computable functions $f$ have a computable inversion. This is because each $f^{-1}(y)$ is representable as the set consisting of the unique path through a $y$-computable pruned tree. In general:
Theorem 3.1.
Proof..
Let $\hat{f}$ be a computable representation of $f$. By compactness
$\forall k\ \ \exists n_{0}\ \ \forall n>n_{0}\ \ \forall\sigma\in 2^{n}:\ |{\hat{f}(\sigma)}|>k.$ (1)
Let $A_{\tau}=\{\sigma:\hat{f}(\sigma)\succeq\tau\}$, $B_{\tau}$ be the $\preceq$-downward closure of $A$,
$B_{\tau}^{n}:=B_{\tau}\cap 2^{n}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptp_{\tau}:=\max\{\hskip 0.85355pt{n}:{|{B_{\tau}^{n}}|=1}\hskip 0.85355pt\}.$
We define a computable representation $\hat{g}$ of the required $g$ by letting $\hat{g}(\tau)$ to be the unique string in $B^{p_{\tau}}_{\tau}$. Then $p_{\tau}$ is nondecreasing, $\hat{g}$ is $\preceq$-monotone and by (1) $B_{\tau}^{n}$, $p_{\tau}$ and $\hat{g}$ are computable. So $\hat{g}$ is a computable representation of a partial computable $g:\subseteq 2^{\omega}\to 2^{\omega}$.
If $\lim_{\tau\prec y}p_{\tau}=m<\infty$ then there are distinct $\sigma,\tau\in 2^{m}$ such that there are arbitrarily long suffixes $\rho$ of both $\sigma$ and $\tau$ with $f(\rho)\prec y$, so $|{f^{-1}(y)}|>1$ by compactness and continuity. Hence
$|{f^{-1}(y)}|=1\ \Rightarrow\ \lim_{\tau\prec y}p_{\tau}=\infty$
which means that $\lim_{n}|{\hat{g}(y\upharpoonright_{n})}|=\infty$ and $g(y)\downarrow$ for $y\in E$. By the definition of $\hat{g}$ we also have $\forall\tau\prec y,\ \hat{f}(\hat{g}(\tau))\preceq\tau$ so $f(g(y))=y$ for $y\in E$. ∎
Corollary 3.2.
Continuously inverting computable non-injective functions is usually hard as they require the commitment of a neighborhood in $f(2^{\omega})$ into an area in the domain of $f$, potentially forbidding the inversion of $f$ on a different neighborhood of $f(2^{\omega})$. We show how to produce such hardness in a simple modification of the zero function by encoding the truth of a $\Sigma^{0}_{1}$ statement into all continuous inversions of a computable $f:2^{\omega}\to 2^{\omega}$.
Example. Let $\exists s,\ E(s)$ be a $\Sigma^{0}_{1}$ formula where $E(s)$ is a computable predicate where $E(s)$ holds for at most one $s$.
We show that any inversion of the computable
$f(x):=\begin{cases}0^{s}\ x(0)\ 0^{\omega}&\textrm{if $E(s)$;}\\ 0^{\omega}&\textrm{otherwise.}\end{cases}$
encodes the answer to whether $\exists s,\ E(s)$ holds.
Let $g$ be a continuous inversion of $f$ and $x:=g(0^{\omega})$. Since $0^{\omega}\in f(2^{\omega})$:
@@ 3 Deterministic inversions
    $x$ is defined and $\exists u,\ g(\llbracket 0^{u}\rrbracket)\subseteq\llbracket x(0)\rrbracket$
@@
    $E(s)$ holds iff $0^{s}10^{\omega}\in f(2^{\omega})$ iff $g(0^{s}10^{\omega})\downarrow$.
We show that $u$ is an upper bound on the $s$ that can satisfy $E(s)$.
By (b) and the continuity of $g$, if $E(s)$ holds:
  • $\hat{g}$ would have to map $0^{s}10^{t}$ into $\llbracket 1\rrbracket$ for some $t$
  • this cannot be done if $\hat{g}$ has already mapped $0^{s+1}$ into $\llbracket 0\rrbracket$.
If $x(0)=0$, by (a) we have $\forall s>u,\ \neg E(s)$
If $x(0)=1$, since $f(x)=f(g(0^{\omega}))=0^{\omega}$ we have $\forall s,\ \neg E(s)$.
[…]
Let $\langle\cdot,\cdot\rangle:\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ be a computable bijection with $\langle n,s\rangle\geq s$ and let $(\emptyset^{\prime}_{s})$ be an effective enumeration of $\emptyset^{\prime}$ without repetitions. We show:
Theorem 3.3.
@@
$f(x;\langle n,s\rangle):=\begin{cases}x(n)&n\in\emptyset^{\prime}_{s+1}-\emptyset^{\prime}_{s}\\ 0&\rm{otherwise.}\end{cases}$
@@
Since $y:=0^{\omega}$ is in the range of $f$, the real $x:=g(y)$ is defined. Let $u_{n}$ be the oracle-use in the computation of $g(y;n)$. It remains to show that
@@
    $x(n)=1\ \Rightarrow\ n\not\in\emptyset^{\prime}$
@@
For (b) assume that $x(n)=0$ and for a contradiction let $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$ for some $s>u_{n}$. Then there exists $z\in 2^{\omega}$ with
$z=f(0^{n-1}1\ast 0^{\omega})=0^{\langle n,s\rangle}10^{\omega}.$
Since $u_{n}$ is the oracle-use of $g(y;n)\downarrow$ and $\langle n,s\rangle\geq s$ we get $0^{u_{n}}\prec z$ and:
@@
4 Randomized inversions
A computable function may be invertible probabilistically by $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ where the secondary argument is viewed as access to a randomness source with respect to $\mu$. Then ‘$g$ inverts $f$’ is the probabilistic event that $g(y,z)\downarrow$, $f(g(y,z))\downarrow=y$ on randomly chosen $y\in f(2^{\omega})$, $z\in 2^{\omega}$.
According to Levin [2] a partial computable $f:\subseteq 2^{\omega}\to 2^{\omega}$ is one-way if no partial computable $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ inverts $f$ with positive probability. The case $\mu(f(2^{\omega}))=0$ is trivial so it is excluded.
Definition 4.1 (Levin [2]).
Given $f:\subseteq 2^{\omega}\to 2^{\omega}$, $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ let
$L_{f}(g):=\{\hskip 0.85355pt{(y,r)}:{g(y,r)\downarrow\ \wedge\ \ f(g(y,r))\downarrow=y}\hskip 0.85355pt\}\subseteq 2^{\omega}\times 2^{\omega}.$
Given $f,g$ the probability that $g$ inverts $f$ refers to $\mu(L_{f}(g))$. If $f$ is partial computable we say that $g$ is a randomized inversion of $f$ if $\mu(L_{f}(g))>0$. We say that $f$ is one-way if $\mu(f(2^{\omega}))>0$ and it does not have any partial computable randomized inversion.
Following Levin [2], instead of $\mu(f(2^{\omega}))>0$ we can require the stronger condition that one-way functions $f$ are random-preserving:
Definition 4.2.
We define random-preserving surjections based on:
@@ 4 Randomized inversions
$f^{-1}(\llbracket V_{i}\rrbracket)=\bigcup_{\tau\in V_{i}}f^{-1}(\llbracket\tau\rrbracket)$
@@
4.1 Construction of a one-way function
We build on Theorem 3.3 using the Lebesgue density theorem ([4, Theorem 1.2.3]). The density of $x\in 2^{\omega}$ in a measurable $G\subseteq 2^{\omega}$ is
$\lim_{n\to\infty}\frac{\mu(\llbracket x\upharpoonright_{n}\rrbracket\cap G)}{\mu(\llbracket x\upharpoonright_{n}\rrbracket)}$
assuming that the limit exists. Lebesgue’s density theorem states that the $\mu$-measure of the $x$ with density $1$ in $G$ is $\mu(G)$.
@@ 4.1 Construction of a one-way function
There is a total computable surjective random-preserving one-way function.
@@
$p(\langle n,s\rangle):=\begin{cases}2n&\textrm{if $n\in\emptyset^{\prime}_{s+1}-\emptyset^{\prime}_{s}$}\\[3.69899pt] 2\langle n,s\rangle+1&\rm{otherwise}\end{cases}$
@@
$f(x;\langle n,s\rangle):=\begin{cases}x(2n)&\textrm{if $n\in\emptyset^{\prime}_{s+1}-\emptyset^{\prime}_{s}$}\\[3.69899pt] x(2\langle n,s\rangle+1)&\rm{otherwise.}\end{cases}$
Since $p$ is a computable injection, Lemma 4.3 implies that $f$ is a total computable random-preserving surjection. It remains to show that $f$ is one-way. Given partial computable $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ and
$L:=\{\hskip 0.85355pt{y\oplus r}:{f(g(y,r))=y}\hskip 0.85355pt\},$
to show that $g$ is not a randomized inversion of $f$ it suffices to show that:
$\big{(}{y\oplus r\in L\ \textrm{ and $y\oplus r$ has density $1$ in $L$}}\big{)}\ \Rightarrow\ \emptyset^{\prime}\leq_{T}y\oplus r.$ (2)
Indeed by [3, 18] we know that $\mu(\{\hskip 0.85355pt{z}:{\emptyset^{\prime}\leq_{T}z}\hskip 0.85355pt\})=0$, so the Lebesgue density theorem would then imply $\mu(L)=\mu(L_{f}(g))=0$.
Assuming the left-hand-side of (2) we fix $k$ such that
$t\geq k\ \Rightarrow\ \frac{\mu(\llbracket(y\oplus r)\upharpoonright_{2t})\rrbracket)\cap L)}{\mu(\llbracket(y\oplus r)\upharpoonright_{2t})\rrbracket)}>\frac{1}{2}$ (3)
and it remains to show $\emptyset^{\prime}\leq_{T}y\oplus r$.
To decide if $n\in\emptyset^{\prime}$, since $g$ is partial computable and $g(y,r)\downarrow$, we can compute $g(y,r;2n)$ and the oracle-use $u$ of this computation. We show:
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,k\}}.$ (4)
For a contradiction assume that $n\in\emptyset^{\prime}_{s+1}-\emptyset^{\prime}_{s}$ for some $s>\max\{u,k\}$.
By the definition of $f$, $L$ and $u$ respectively, for all $z,w$ we have
  1.    (a)
    $f(g(z,w);\langle n,s\rangle)=g(z,w;p(\langle n,s\rangle))=g(z,w;2n)$
  2.    (b)
    $z\oplus w\in L\ \Rightarrow\ f(g(z,w);\langle n,s\rangle)=z(\langle n,s\rangle)$
  3.    (c)
    $(z,w)\in\llbracket y\upharpoonright_{s}\rrbracket\times\llbracket r\upharpoonright_{s}\rrbracket\ \Rightarrow\ g(z,w;2n)=g(y,r;2n)$.
Combining (a), (b), (c) we get
$z(\langle n,s\rangle)=g(y,r;2n)\hskip 8.5359pt\textrm{for all $z\oplus w\in\llbracket(y\oplus r)\upharpoonright_{2s}\rrbracket\cap L$}$
so all reals in $\llbracket(y\oplus r)\upharpoonright_{2s}\rrbracket\cap L$ agree on position $2\langle n,s\rangle$. This contradicts (3) since $\langle n,s\rangle\geq s>k$. So (4) holds, which completes the proof of (2). ∎

### v3

# George Barmpalias State Key Lab of Computer Science

@@ 1 Introduction
A function is one-way if it is computationally easy to compute but hard to invert, even probabilistically. Many aspects of computer science such as computational complexity, pseudorandom generators and digital signatures rely on one-way functions from strings (finite binary sequences) to strings [20, 19]. Although their existence is not known, much research effort has focused on their implications and relations with other fundamental problems in computation [15, 14, 11]. Recent research has established strong connections between one-way functions and Kolmogorov complexity [26, 21, 22, 13].
Levin [20, §2.2] asked for the existence of a one-way function from reals (infinite binary sequences) to reals, which he defined in terms of probabilistic computability with respect to the uniform measure $\mu$ [^1] This was also discussed in [6]. These are partial computable functions $f$ which preserve algorithmic randomness and
the probability that $M$ inverts $f$, namely $f(M(y))=y$, is zero
for each probabilistic Turing machine $M$. We give a positive answer:
@@
There exists a total computable one-way surjection $f$.
The hardness of inverting $f$ is based on the hardness of the halting problem $\emptyset^{\prime}$. As a result, every randomized continuous inversion of $f$ computes $\emptyset^{\prime}$.
Gács [9] independently constructed $f$ which is ‘one-way’ with respect to the domain instead of the range of $f$: for each partial computable $g$,
$\mu\big{(}{\{\hskip 0.85355pt{(x,r)}:{f(g(f(x),r))=f(x)}\hskip 0.85355pt\}}\big{)}=0$ (1)
and $f$ is partial computable with domain of positive uniform measure $\mu$.
Although (1) appears to be a closer analogue to the one-way functions in computational complexity, it does not preserve randomness, which is required by Levin [20]. Effective partial maps that meet (1) often have a null image. Levin’s definition implies (1) while the converse fails [5, §3.2].
In §3 we justify Levin’s definition as the correct analogue of one-way functions on discrete domains, over weaker formalizations. We exhibit a one-way real function using a framework that can be adapted toward meeting additional conditions, and explore their properties.
We do not know if there is an injective one-way real function. In §4 we explore the extent to which injectivity is compatible with one-way real functions and end with a summary and some problems in §5.
@@ 2 Preliminaries
Let $\mathbb{N}$ be the set of natural numbers, represented by $n,m,i,j,t,s$. Let
  • $2^{\omega}$ be the set of reals, represented by variables $x,y,z,v,w$
  • $2^{<\omega}$ the set of strings, represented by variables $\sigma,\tau,\rho$.
We index the bits $x(i)$ of $x$ starting from $i=0$. Let
  • $x\upharpoonright_{n}$ be the $n$-bit prefix $x(0)x(1)\cdots x(n-1)$ of $x$
  • $\preceq$, $\prec$ be the prefix and strict prefix relation between strings
  • $\sigma\mid\tau$ denotes that $\sigma\not\preceq\tau\ \wedge\ \tau\not\preceq\sigma$
and $\succeq$,$\succ$ , ≻ denote the suffix relations which, along with the prefix relations can also apply between a string and a real. Given $x,y$ let
@@ 2.1 Computability and randomness
which we call cylinders. Let $\mu$ be the uniform measure on $2^{\omega}$, determined by $\mu(\llbracket\sigma\rrbracket)=2^{-|\sigma|}$. Probability in $2^{\omega}\times 2^{\omega}$ is reduced to $2^{\omega}$ via the measure-preserving $(x,y)\mapsto x\oplus y$. Classes of $\mu$-measure 0 are called null.
A tree is a downward $\preceq$-closed $T\subseteq 2^{<\omega}$ and
  • $T$ is pruned if every $\sigma\in T$ has an extension in $T$
  • $T$ is perfect if every $\sigma\in T$ has two extensions $\tau|\rho$ in $T$.
  • $x$ is a path through $T$ if all of its prefixes belong to $T$.
Let $[T]$ be the class of all paths through $T$.
We use the standard notion of relative computability in terms of Turing machines with oracles from $2^{\omega}$. Turing reducibility $x\leq_{T}z$ means that $x$ is computable from $z$ (is $z$-computable) and is a preorder calibrating $2^{\omega}$ according to computational power in the Turing degrees.
Effectively open sets, also known as $\Sigma^{0}_{1}$ classes, are subsets of $2^{\omega}$ of the form
$\bigcup_{i}\llbracket\sigma_{i}\rrbracket\hskip 8.5359pt\textrm{where $(\sigma_{i})$ is computable.}$.
Effectively closed sets or $\Pi^{0}_{1}$ classes are the complements of $\Sigma^{0}_{1}$ classes. Every $\Pi^{0}_{1}$ class is the set of paths through some computable tree, and vice versa.
If $(\sigma_{n,i})$ is computable then
@@
Equivalently, $V\in\Pi^{0}_{2}$ iff there is a computable predicate $P$ such that
$x\in V\iff\forall n\ \exists s\ P(x\upharpoonright_{n},x\upharpoonright_{s}).$
Relativization to an oracle $r$ defines $\Sigma^{0}_{1}(r),\Pi^{0}_{2}(r)$ classes and so on.
A Martin-Löf test is a uniformly $\Sigma^{0}_{1}$ sequence $(V_{n})$ with $\mu(V_{n})\leq 2^{-n}$.
Definition 2.1.
Given $k\in\mathbb{N}$, a real $x$ is
  • random if $x\notin\bigcap_{n}V_{n}$ for any Martin-Löf test $(V_{n})$
  • weakly random if it is in every $\Sigma^{0}_{1}$ set of measure $1$.
  • weakly $k$-random if it is in every $\Sigma^{0}_{k}$ set of measure $1$.
Relativization to oracle $r$ defines $r$-random, weakly $r$-random and so on.
By the countable additivity of the uniform measure $\mu$:
$x$ is weakly $r$-random iff it is not in any null $\Sigma^{0}_{2}(r)$ class (2)
and similarly for weakly $k$-random reals $x$.
2.2 Computable analysis
A Turing machine with a one-way infinite output tape and access to an oracle from $2^{\omega}$ may eventually print an infinite binary sequence, hence defining a partial map from reals to reals. This standard notion of computability of real functions [25] is almost as old as computability itself [10] and implies that computable real functions are continuous. Let
@@ 2.2 Computable analysis
  • $f(x)\downarrow,f(x)\uparrow$ denote that $f(x)$ is defined or undefined
  • $f(x;n):=f(x)(n)$ denote bit $n$ of $f(x)$
and $f(x;n)\downarrow,f(x;n)\uparrow$ denote that $f(x;n)$ is defined or undefined.
If $f$ is partial computable, the oracle-use of $f(x;n)$ is the prefix of $x$ that has been read by the underlying Turing machine at the time where when $f(x;n)$ is printed on position $n$ of the one-way output tape.
Definition 2.2.
The range and the inverse images of $f$ are denoted by
$f(2^{\omega}):=\{\hskip 0.85355pt{y}:{\exists x,\ f(x)=y}\hskip 0.85355pt\}\hskip 8.5359pt\textrm{and}\hskip 8.5359ptf^{-1}(y):=\{\hskip 0.85355pt{x}:{f(x)=y}\hskip 0.85355pt\}.$
If $f$ is total and continuous the inverse images
@@
are closed and effectively closed if $f$ is also computable. Continuous functions $f:\subseteq 2^{\omega}\to 2^{\omega}$ can be defined via a representation: a $\subseteq$-monotone map between cylinders. Formally, let $\lim_{\tau\prec x}\hat{f}(\tau)=z$ denote that
[…]
Given a continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ we say that $g:\subseteq 2^{\omega}\to 2^{\omega}$:
  • inverts $f$ on $y$ if $g(y)\downarrow$ and $f(g(y))\downarrow=y$
We say that $\hat{f}:2^{<\omega}\to 2^{<\omega}$ is a representation of $f$ if
  • $\sigma\preceq\tau\implies\hat{f}(\sigma)\preceq\hat{f}(\tau)$
  • $f(x)\downarrow\iff\lim_{\tau\prec x}\hat{f}(\tau)=f(x)\iff\lim_{\tau\prec x}|{\hat{f}(\tau)}|=\infty$.
Every continuous $f:\subseteq 2^{\omega}\to 2^{\omega}$ has a representation. Every partial computable $f$ has a computable representation. Under an effective coding of the graphs of representations into $2^{\omega}$ we identify them with their codes.
@@
Note that by [24] there is no canonical way to assign a representation to each continuous $f$: there are continuous $f$ such that for each representation $z$ of $f$ there is another one of lesser Turing degree.
3 One-way functions
Toward foundational research that is relevant to mathematical practice, Levin proposed a new direction [20]. [^2] This includes 15 drafts developed from 2022 to the end of 2024. He used invertibility of real functions to express the axioms of choice and powerset, and emphasized the significance of its computational hardness in the context of his proposal.
He formally extends the notion of one-way functions to continuous domains and he asks about their existence. One may formalize this notion in several ways which are not equivalent, especially in the case of partial computable maps. In §3.1 we give the formal definition from [20] and explain why it is the correct extension of the discrete notion over weaker alternatives.
In §3.2 we construct a total computable one-way surjection. We do not know whether there is a one-way injection on the reals. In §3.3 we establish properties of one-way real functions relating to this question, setting the basis for the comprehensive analysis in §4.
3.1 Levin’s one-way functions on the reals
In computational complexity a function from strings to strings is one-way if it is easy to compute and hard to invert, even probabilistically.
By Levin [20] a one-way real function $f$ must
@@ 3.1 Levin’s one-way functions on the reals
    be partial computable with domain of positive measure
@@
    have no effective probabilistic inversion
  3.    (c)
    satisfy $\mu(f^{-1}(\llbracket\tau\rrbracket))=\textsf{O}\hskip 0.28436pt\big{(}{\mu(\llbracket\tau\rrbracket)}\big{)}$.
Conditions (a), (b) clearly correspond to conditions in the discrete one-way functions. Probabilistic inversions of $f$ are facilitated by $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ where the secondary argument represents access to a randomness source. However probability in (b) may refer to the domain or the range of $f$.
Levin [20] chose the latter, interpreting “$g$ inverts $f$” as the event that
$f(g(y,r))\downarrow=y$ on randomly chosen $y\in f(2^{\omega})$, $r\in 2^{\omega}$. (3)
Formally, the probability that $g$ inverts $f$ is the measure of
$L_{f}(g):=\{\hskip 0.85355pt{y\oplus r}:{g(y,r)\downarrow\ \wedge\ \ f(g(y,r))\downarrow=y}\hskip 0.85355pt\}$
and $g$ is a randomized inversion of $f$ if $\mu(L_{f}(g))>0$.
Definition 3.1 (Levin [20]).
We say that $f:\subseteq 2^{\omega}\to 2^{\omega}$ is one-way if it
  • is partial computable, random-preserving with positive domain
  • does not have any partial computable randomized inversion.
The alternative is condition (1) of Gács [9] which, as discussed in §1, is strictly weaker if we assume that $f$ is random-preserving.
The mild measure-preservation condition (c) is equivalent to randomness-preservation and implies $\mu(f(2^{\omega}))>0$ for partial computable $f$ (see [5, §3.2]). To see why (c) is essential consider the weaker conditions:
  1. (i)
    $\mu(f(E))>0$ for each subset $E$ of the domain of $f$ with $\mu(E)>0$
  2. (ii)
    the domain of $f$ contains $R$ with $\mu(R)>0$ and $\mu(f(R))>0$
  3. (iii)
    there is a $\emptyset^{\prime}$-random $x$ such that $f(x)$ is random
and note that (c)$\to$(i)$\to$(ii)$\to$(iii) while none of the converses holds. It is not hard to show (see [5, §3.2]) that each (i), (ii), (iii) is equivalent to (c) up to effective restrictions of $f$. So essentially (c) only asks that $f$ maps at least one sufficiently random real to a random real.
Weaker versions based on combinations of (i), (ii), (iii), $\mu(f(2^{\omega}))>0$ are less robust as they do not form a discernible hierarchy [5].
3.2 Construction of a total one-way surjection
We will map input $x$ to a permutation $y$ of selected bits of $x$. Let
  • $\langle\cdot,\cdot\rangle:\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ be a computable bijection with $\langle n,s\rangle\geq s$
  • $(\emptyset^{\prime}_{s})$ be an effective enumeration of $\emptyset^{\prime}$ without repetitions.
We give a simplified construction which we extend in §3.2 for the full result.
Proposition 3.2.
@@ 3.2 Construction of a total one-way surjection
[…]
@@
Since $y:=0^{\omega}$ is in the range of $f$, the real $x:=g(y)$ is defined. Let $u_{n}$ be the oracle-use in the computation of $g(y;n)$, that is, the minimum $u$ such that the computation of $g(y;n)$ reads only $y\upharpoonright_{u}$. It remains to show that
@@
    if $x(n)=1$ then $n\not\in\emptyset^{\prime}$
@@
For (b) assume that $x(n)=0$ and for a contradiction let $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$ for some $s>u_{n}$. Then there is $z\in 2^{\omega}$ with $z=f(0^{n}10^{\omega})=0^{\langle n,s\rangle}10^{\omega}$.
Since $u_{n}$ is the oracle-use of $g(y;n)\downarrow$ and $\langle n,s\rangle\geq s$ we get $0^{u_{n}}\prec z$ and
@@
Our one-way function will be a permutation of selected bits of the input. We show that such maps meet certain properties required of one-way functions.
@@
$f^{-1}(\llbracket V_{i}\rrbracket)=\bigcup_{\tau\in V_{i}}f^{-1}(\llbracket\tau\rrbracket).$
@@
We extend the argument in Proposition 3.2 to obtain a one-way function.
@@
There is a total computable random-preserving one-way surjection, such that every randomized inversion of it computes $\emptyset^{\prime}$.
@@
$p(\langle n,s\rangle):=\begin{cases}\ \ \ \ 2n&\textrm{if $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$}\\[3.69899pt] \ 2\langle n,s\rangle+1&\rm{otherwise}\end{cases}$
@@
$f(x;\langle n,s\rangle):=\begin{cases}\ \ \ \ x(2n)&\textrm{if $n\in\emptyset^{\prime}_{s}-\emptyset^{\prime}_{s-1}$}\\[3.69899pt] \ x(2\langle n,s\rangle+1)&\rm{otherwise.}\end{cases}$
Since $p$ is a computable injection, Lemma 3.3 implies that $f$ is a total computable random-preserving surjection. Given $g:\subseteq 2^{\omega}\times 2^{\omega}\to 2^{\omega}$ and